OpenStudy (anonymous):

A discriminant of 0 indicates what type of roots? Answer Two Rational Two Complex One Rational Double Two Irrational

OpenStudy (jamesj):

Well, look at an example: x^2 - 2x + 1 The discriminant = 2^2 - 4(1)(1) = 4 - 4 =0 and x^2 - 2x + 1 = (x-1)^2 => x = 1 is a double root So it looks like discriminant = 0 means .... what?